# Waveform Design of Multi-User-Multi-Target ISAC System based on Kullback–Leibler Divergence

### v1

@@ VII Numerical Results
In this section, we present the simulation results for both our conventional benchmarks and the two optimisation techniques introduced previously. The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout this section. The total number of antennas at BS is $N=20$, a number of $L=100$ snapshots, the maximum number of potential radar targets is $T=3$, and the number of UEs is $K=3$. The pathloss exponent is $\eta=3$ to model the effect of large-scale fading, and the distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters, respectively, and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters, respectively. The channel variance is $\sigma_{h}^{2}=1$.
@@ VII-A Trade-off for the conventional benchmarks
The conventional identity covariance matrix design is used for the radar subsystem with a covariance matrix is set to $\mathbf{R}_{w}=\mathbf{I}_{N}$, whereas the communication subsystem uses MRT and ZF precoding introduced earlier.
@@
Figure 2: The performance of the communication subsystem for both beamforming schemes, where the Bit Error Rate (BER) and the achiveable $\mathrm{KLD}_{\mathrm{c}}$ are shown versus $P_{\mathrm{R}}/{N}_{o}$, with $P_{\mathrm{r}}=0.5$.
@@
Fig. 2 presents the performance of the communication subsystem for both ZF and MRT beamforming schemes. In Fig. 2.a, the MRT scheme’s performance is depicted, with the KLD performance on the right $y$-axis and the BER performance on the left. As can be observed from the figure, the KLD values for all UEs increase with $P_{\mathrm{R}}/{N}_{o}$ until around $10$ dB, where a noticeable upper-bound occurs primarily caused by radar interference, with IUI contributing to a less extent. Similarly, the BER decreases with $P_{\mathrm{R}}/{N}_{o}$, indicating an improvement in the communication subsystem performance, but also exhibits the same error floor at around $10$ dB. Fig. 2.b illustrates the ZF scheme’s performance, exhibiting similar trends in the KLD and BER curves. However, compared to the MRT scheme, the ZF scheme shows an increase in KLD values and a decrease in BER values for the same received SNR, indicating that the primary source of interference is indeed caused by radar interference as IUI is cancelled using ZF precoding. Consequently, both the ZF and MRT schemes exhibit a drastic error floor in BER, or equivalently an upper bound for KLD, due to the unmitigated radar interference at $P_{\mathrm{c}}=0.5$. Moreover, this figure clearly shows the effectiveness of using KLD to characterize the communication system.
@@
Figure 3: The receiver operating characteristic (ROC) curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios.
@@
Fig. 3 shows the receiver operating characteristic (ROC) curves depicting the performance of the radar subsystem for the three potential targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios, for the same ISAC system in Fig. 2 with ZF precoding for communication subsystem assumed. The ROC curves are obtained by plotting the probability of detection ($P_{D}$) versus the probability of false alarm ($P_{FA}$) for each target. In Fig. 3, where $P_{\mathrm{R}}/{N}_{o}=0$ dB, the ROC curves for all three targets $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ are relatively close to each other, indicating similar detection performance at low signal-to-noise ratios. In Fig. 3, with $P_{\mathrm{R}}/{N}_{o}=10$ dB, the ROC curves for the three targets start to diverge, where $\mathrm{T}_{3}$ exhibits the best performance, followed by $\mathrm{T}_{1}$ followed by $\mathrm{T}_{2}$. This suggests that as the signal-to-noise ratio increases, the radar subsystem’s ability to detect the targets and distinguish between them improves. Finally, in Fig. 3, in which $P_{\mathrm{R}}/{N}_{o}=20$ dB, the separation between the ROC curves becomes even more pronounced. As can be observed in 3, at $P_{FA}=10^{-3}$, $P_{D}$ is $\{0.909,0.804,0.935\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, where $\mathrm{T}_{3}$ maintains its superior performance, while $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$ show significant improvements in their detection probabilities compared to the lower $P_{\mathrm{R}}/{N}_{o}$ scenarios. This demonstrates that at high signal-to-noise ratios, the radar subsystem can effectively detect and differentiate between the three targets, with $\mathrm{T}_{3}$ being the most easily detectable as it is the closest target to the BS.
@@
Figure 4: a) The performance of the radar subsystem where the achievable $\mathrm{KLD_{\mathrm{r}}}$ is shown versus $P_{\mathrm{R}}/{N}_{o}$, across different values of $P_{\mathrm{r}}$. b) The KLD performance trade-off between the radar and communication subsystems for ZF and MRT
In Fig.4.a, the performance of the radar subsystem is illustrated through the achievable $\mathrm{KLD_{\mathrm{r}}}$ of each target versus the $P_{\mathrm{R}}/{N}_{o}$ across different values of $P_{\mathrm{r}}$. It is shown that as $P_{\mathrm{R}}/{N}_{o}$ increases, the achievable $\mathrm{KLD_{\mathrm{r}}}$ for each target also increases for all values of $P_{\mathrm{r}}$. This indicates that the radar subsystem’s performance improves with higher $P_{\mathrm{R}}/{N}_{o}$. The performance curves for different $P_{\mathrm{r}}$ values are relatively close to each other, suggesting that the system’s performance is less sensitive to changes in $P_{\mathrm{r}}$ because the communication subsystem signal is also incorporated in the target detection, using our improved system model.
@@
Fig. 4.b presents the trade-off performance between radar and communication subsystems using the ZF and MRT for the communication subsystem and the conventional identity covariance design for the radar subsystem. The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ at equal power allocation between both subsystems, with each point on the curve representing a different $P_{\mathrm{R}}/N_{o}$ value. The $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ outperforms the $\mathrm{KLD_{c}}^{\mathrm{MRT}}$ slightly, as both of them exhibits saturation at higher $P_{\mathrm{R}}/N_{o}$ due to the destructive interference from the radar subsystem mainly, while the $\mathrm{KLD_{r}}$ shows a steady increase when moving from left to right on the curve. It can be observed from the figure that a $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ of 8 bits can be achieved at $P_{\mathrm{R}}/N_{o}=12$ dB for $\mathrm{UE_{2}}$, whereas the same amount of $\mathrm{KLD_{c}}^{\mathrm{MRT}}$ is achieved at around $25$ dB, which implies an SNR gain of about $13$ dB in favour of ZF.
@@ VII-B Radar Waveform KLD-based Optimisation
The radar waveform KLD-based optimization algorithm is configured with a maximum iteration limit of $\mathrm{max\_iter}=1000$ and a convergence tolerance of $\epsilon=10^{-6}$. The penalty parameter is initialized at $\rho_{0}=1$, with a penalty increase parameter of $\gamma=1.5$, and an initial step size of $\alpha_{0}=0.1$. The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are set to 10 bits.
@@
Fig. 5 illustrates the performance of the radar and communication subsystems using the optimized radar waveform. Fig. 5.a shows the achievable $\mathrm{KLD_{r}}$ for each target as a function of $P_{\mathrm{R}}/N_{o}$. As can be seen from the figure, as $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves for all targets, indicating enhanced radar performance at higher power-to-noise ratios. It can be also observed that target T3 consistently achieves the highest $\mathrm{KLD_{r}}$, followed by T1 and T2, reflecting their respective distances from the BS. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits for T1, T2, and T3 respectively. These values increase to about $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB, showing improvement in radar performance. The rate of $\mathrm{KLD_{r}}$ improvement after 15 dB appears to enter a near-linear relationship with $P_{\mathrm{R}}/N_{o}$, suggesting a consistent pattern of performance enhancement in the radar system.
Fig. 5.b depicts the communication subsystem performance through $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$. The $\mathrm{KLD_{c}}$ for $P_{\mathrm{R}}/N_{o}$ values between 00-$8$ dB exhibits a minimal improvement, indicative of system infeasibility in this low power regime. This minimal growth at the low SNR region is attributed to the system’s inability to meet the minimum KLD requirements for all UEs simultaneously. UE3 consistently achieves the highest $\mathrm{KLD_{c}}$, followed by UE1 and UE2, which aligns with their respective distances from the BS. At $P_{\mathrm{R}}/N_{o}=10$ dB, $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ values are approximately $\{55.14,20.09,186.09\}$ for UE1, UE2, and UE3 respectively. These values increase substantially to about $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The sharp increase in $\mathrm{KLD_{c}}$ beyond $8$ dB for all UEs suggests that at this point, the system has reached its optimization feasibility point. Subsequently, the communication performance is enhanced significantly at higher SNR and after reaching feasibility as the radar interference is effectively reduced. These results demonstrate the intricate relationship between radar detection capabilities, and communication system performance in this dual-function system. The analysis shows how the system transitions from infeasibility to high performance as power increases, with both subsystems benefiting once the feasibility threshold is surpassed. The consistent ordering of performance among targets and UEs reflects the impact of their respective distances from the BS on system performance.
@@
Fig. 6.a shows the receiver operating characteristic (ROC) curves depicting the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. At $P_{\mathrm{R}}/{N}_{o}=20$ dB, and $P_{FA}=10^{-3}$, the $P_{D}$ is $\{0.997,0.991,0.998\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, while for the non-optimal way the $P_{D}$ is $\{0.909,0.804,0.935\}$. This demonstrates that the optimisation improves detection capabilities comparing it with non-optimisation in Fig. 3. Fig. 6.b depicts the BER performance for three UEs versus $P_{\mathrm{R}}/N_{o}$. The BER decreases as $P_{\mathrm{R}}/N_{o}$ increases for all three UEs, indicating improved communication performance at higher $P_{\mathrm{R}}/N_{o}$. $\mathrm{UE}_{3}$ consistently achieves the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. The performance of the communication subsystem improves significantly with radar waveform optimization. In the non-optimal case, $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB, as shown in Fig. 2. However, with optimization, $P_{e}$ drops below $10^{-5}$ by $13$ dB across all UEs, due to the communication subsystem’s sensitivity to radar interference, which is better mitigated through optimized waveforms.
@@
Fig. 7.a illustrates the trade-off performance between radar and communication subsystems using optimized radar waveform. The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$, with each point on the curve representing a different $P_{\mathrm{R}}/N_{o}$ value. Moving from left to right, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ increase, indicating that higher $P_{\mathrm{R}}/N_{o}$ values lead to improved performance in both subsystems. This visualization provides insights into system behaviour across different $P_{\mathrm{R}}/N_{o}$ levels, highlighting the optimization technique’s robustness in balancing radar and communication performance, we can observe that the optimal way shows an increasing and favourable trade-off, by comparison to the non optimal way in Fig. 4.b, where the $\mathrm{KLD_{c}}$ saturates at around $11$ bits as $P_{\mathrm{R}}/N_{o}$ increases.
Fig. 7.b shows the performance of the radar waveform optimization technique in terms of average CPU time versus $P_{\mathrm{R}}/N_{o}$. The plot reveals that the computational complexity of the optimization algorithm remains relatively stable across different $P_{\mathrm{R}}/N_{o}$ values. The average CPU time fluctuates slightly between $0.03$ and $0.035$ seconds, with no clear trend of increase or decrease as $P_{\mathrm{R}}/N_{o}$ changes. This suggests that the optimization technique maintains consistent computational efficiency regardless of the $P_{\mathrm{R}}/N_{o}$ values, which is advantageous for practical implementation in varying channel conditions.

### v2

@@ VII Numerical Results
In this section, we present simulation results for our conventional benchmarks and the two previously introduced optimisation techniques. The total transmit power is fixed at $P_{\mathrm{T}}=1$, and QPSK modulation is used throughout. The base station has $N=20$ antennas, with $L=100$ snapshots, a maximum of $T=3$ potential radar targets, and $K=3$ UEs. The pathloss exponent is $\eta=3$ to model large-scale fading. The distances from the BS to the UEs are $d_{\mathrm{c},k}=\{150,210,100\}\forall k\leq K$ meters and to the targets are $d_{\mathrm{r},t}=\{100,115,95\}\forall t\leq T$ meters. The channel variance is $\sigma_{h}^{2}=1$.
@@ VII-A Trade-off for the conventional benchmarks
The radar subsystem employs a conventional identity covariance matrix design, with $\mathbf{R}_{w}=\mathbf{I}_{N}$, while the communication subsystem utilises the previously introduced MRT and ZF.
@@
Figure 1: Performance of the communication subsystem for both beamforming schemes, showing BER and achievable $\mathrm{KLD}_{\mathrm{c}}$ versus $P_{\mathrm{R}}/{N}_{o}$ with $P_{\mathrm{r}}=0.5$.
@@
Fig. 1 shows the communication subsystem performance for both ZF and MRT beamforming. In Fig. 1.a, the MRT scheme’s performance is illustrated, with KLD on the right $y$-axis and BER on the left. The KLD values for all UEs increase with $P_{\mathrm{R}}/{N}_{o}$ up to about $10$ dB, where an upper bound appears due to radar interference, with IUI contributing less significantly. Similarly, the BER decreases with $P_{\mathrm{R}}/{N}_{o}$, reflecting improved communication performance but also reaches an error floor around $10$ dB. Fig. 1.b presents the ZF scheme’s performance, showing similar KLD and BER trends. However, the ZF scheme exhibits higher KLD values and lower BER values for the same received SNR compared to MRT, confirming that radar interference is the primary source, as IUI is cancelled using ZF precoding. Both schemes display a significant error floor in BER and an upper bound in KLD due to persistent radar interference at $P_{\mathrm{c}}=0.5$. This figure effectively demonstrates the utility of KLD in characterising the communication system.
@@
Figure 2: The ROC curves depict the performance for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios.
@@
Fig. 2 presents the receiver operating characteristic (ROC) curves for the radar subsystem’s performance across three potential targets at varying $P_{\mathrm{R}}/{N}_{o}$ levels, using ZF precoding for the communication subsystem as shown in Fig. 1. The ROC curves plot the probability of detection ($P_{D}$) against the probability of false alarm ($P_{FA}$) for each target. At $P_{\mathrm{R}}/{N}_{o}=0$ dB, the ROC curves for targets $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$ are closely aligned, indicating similar detection performance at low SNR. However, at $P_{\mathrm{R}}/{N}_{o}=10$ dB, the curves diverge, with $\mathrm{T}_{3}$ showing the best performance, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$. This trend continues at $P_{\mathrm{R}}/{N}_{o}=20$ dB, where the separation is even more pronounced. Notably, at $P_{FA}=10^{-3}$, the $P_{D}$ values are $\{0.909,0.804,0.935\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively. $\mathrm{T}_{3}$ remains the most detectable, while both $\mathrm{T}_{2}$ and $\mathrm{T}_{1}$ show significant detection improvements compared to lower $P_{\mathrm{R}}/{N}_{o}$ scenarios. This illustrates that higher SNR enhances the radar subsystem’s capability to detect and differentiate targets effectively.
@@
Figure 3: a) Radar subsystem performance: achievable $\mathrm{KLD_{\mathrm{r}}}$ versus $P_{\mathrm{R}}/{N}_{o}$ for varying $P_{\mathrm{r}}$. b) KLD trade-off between subsystems for ZF and MRT.
@@
In Fig. 3.a, the radar subsystem performance is illustrated through the achievable $\mathrm{KLD_{\mathrm{r}}}$ of each target versus $P_{\mathrm{R}}/{N}_{o}$ for varying $P_{\mathrm{r}}$. As $P_{\mathrm{R}}/{N}_{o}$ increases, the achievable $\mathrm{KLD_{\mathrm{r}}}$ also improves, indicating enhanced radar performance. The performance curves for different $P_{\mathrm{r}}$ values are close, suggesting that the system is less sensitive to $P_{\mathrm{r}}$ changes due to the communication signal’s incorporation in target detection using our improved system model. Fig. 3.b presents the trade-off performance between both subsystems using ZF and MRT for communication, along with the conventional identity covariance design for radar. The plot shows $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ at equal power allocation, with each point representing a different $P_{\mathrm{R}}/N_{o}$ value going higher from left to right. The $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ slightly outperforms $\mathrm{KLD_{c}}^{\mathrm{MRT}}$, both exhibiting saturation at higher $P_{\mathrm{R}}/N_{o}$ due to radar interference, while $\mathrm{KLD_{r}}$ steadily increases along the curve. Notably, a $\mathrm{KLD_{c}}^{\mathrm{ZF}}$ of 8 bits is achieved at $P_{\mathrm{R}}/N_{o}=12$ dB for $\mathrm{UE_{2}}$, whereas the same $\mathrm{KLD_{c}}^{\mathrm{MRT}}$ requires around $25$ dB, implying an SNR gain of about $13$ dB in favor of ZF.
@@ VII-B Radar Waveform KLD-based Optimisation
The radar waveform KLD-based optimisation algorithm is set with a maximum of $\mathrm{max\_iter}=1000$, convergence tolerance $\epsilon=10^{-6}$, initial penalty $\rho_{0}=1$, penalty increase $\gamma=1.5$, and initial step size $\alpha_{0}=0.1$. The KLD lower bounds $\underline{A}_{t}$ and $\underline{B}_{k}$ for all targets and UEs are 10 bits.
@@
Fig. 4 illustrates the performance of both subsystems with the optimised radar waveform. In Fig. 4.a, the achievable $\mathrm{KLD_{r}}$ for each target versus $P_{\mathrm{R}}/N_{o}$ is shown. As $P_{\mathrm{R}}/N_{o}$ increases, $\mathrm{KLD_{r}}$ improves for all targets, indicating enhanced radar performance. $\mathrm{T}_{3}$ consistently achieves the highest $\mathrm{KLD_{r}}$, followed by $\mathrm{T}_{1}$ and $\mathrm{T}_{2}$, reflecting their distances from the BS. At $P_{\mathrm{R}}/N_{o}=10$ dB, the $\{\mathrm{KLD}_{\mathrm{r},1},\mathrm{KLD}_{\mathrm{r},2},\mathrm{KLD}_{\mathrm{r},3}\}$ values are approximately $\{11.78,8.06,13.38\}$ bits for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, increasing to $\{48.06,39.68,51.31\}$ bits at $P_{\mathrm{R}}/N_{o}=20$ dB. The rate of $\mathrm{KLD_{r}}$ improvement after 15 dB suggests a near-linear relationship with $P_{\mathrm{R}}/N_{o}$, indicating consistent performance enhancement in the radar system. Fig. 4.b illustrates the communication subsystem performance, showing $\mathrm{KLD_{c}}$ versus $P_{\mathrm{R}}/N_{o}$. For $P_{\mathrm{R}}/N_{o}$ values between 00-$8$ dB, $\mathrm{KLD_{c}}$ exhibits minimal improvement, indicating system infeasibility in this low power regime due to the inability to meet minimum KLD requirements for all UEs. $\mathrm{UE}_{3}$ consistently achieves the highest $\mathrm{KLD_{c}}$, followed by $\mathrm{UE}_{1}$ and $\mathrm{UE}_{2}$. At $P_{\mathrm{R}}/N_{o}=10$ dB, the values $\{\mathrm{KLD}_{\mathrm{c},1},\mathrm{KLD}_{\mathrm{c},2},\mathrm{KLD}_{\mathrm{c},3}\}$ are approximately $\{55.14,20.09,186.09\}$ for $\mathrm{UE}_{1}$, $\mathrm{UE}_{2}$, and $\mathrm{UE}_{3}$, increasing to $\{551.38,200.94,1860.91\}$ at $P_{\mathrm{R}}/N_{o}=20$ dB. The sharp increase in $\mathrm{KLD_{c}}$ beyond $8$ dB suggests the system has reached its optimisation feasibility point, significantly enhancing the communication performance as the radar interference is effectively reduced. These results highlight the intricate relationship between radar detection capabilities and communication system performance, illustrating how both subsystems benefit once the feasibility threshold is surpassed, with the performance reflecting the respective distances from the BS.
@@
Fig. 5.a presents the ROC curves for the three targets across different $P_{\mathrm{R}}/{N}_{o}$ scenarios. At $P_{\mathrm{R}}/{N}_{o}=20$ dB and $P_{FA}=10^{-3}$, the $P_{D}$ is $\{0.997,0.991,0.998\}$ for $\mathrm{T}_{1}$, $\mathrm{T}_{2}$, and $\mathrm{T}_{3}$, respectively, compared to $\{0.909,0.804,0.935\}$ in the non-optimal case, demonstrating improved detection capabilities through optimisation. Fig. 5.b shows the BER performance for the three UEs versus $P_{\mathrm{R}}/N_{o}$. As $P_{\mathrm{R}}/N_{o}$ increases, BER decreases for all UEs, with $\mathrm{UE}_{3}$ achieving the lowest BER, followed by $\mathrm{UE}_{2}$ and $\mathrm{UE}_{1}$. The communication subsystem significantly benefits from radar waveform optimisation, as evidenced by the non-optimal case where $P_{e}$ saturates at $8\times 10^{-3}$ after 15 dB, as shown in Fig. 1, while optimised waveforms reduce $P_{e}$ below $10^{-5}$ by $13$ dB, effectively mitigating radar interference.
@@
Fig. 6.a illustrates the trade-off performance between both subsystems using the optimised radar waveform, showing $\mathrm{KLD_{c}}$ versus $\mathrm{KLD_{r}}$ for different $P_{\mathrm{R}}/N_{o}$ values. As $P_{\mathrm{R}}/N_{o}$ increases, both $\mathrm{KLD_{c}}$ and $\mathrm{KLD_{r}}$ improve, indicating enhanced performance in both subsystems. This highlights the robustness of the optimisation technique in balancing radar and communication performance. In comparison, the non-optimal approach in Fig. 3.b saturates at around $11$ bits, indicating an increasing and favourable trade-off for the optimal way. Fig. 6.b presents the average CPU time of the radar waveform optimisation technique versus $P_{\mathrm{R}}/N_{o}$. The average CPU time fluctuates slightly between $0.03$ and $0.035$ seconds, with no clear trend of increase or decrease as $P_{\mathrm{R}}/N_{o}$ changes, indicating consistent computational efficiency across varying $P_{\mathrm{R}}/N_{o}$ values, which is beneficial for practical implementation in diverse channel conditions.